OpenStudy (anonymous):

At a used bike store, dirt bikes sell for 1/2 of their original price, and racing bikes sell for 1/3 of their original price. The original prices of dirt bikes range from $25 to $75, and the original prices of racing bikes range from $100 to $150. Which of these is a reasonable total cost to pay at the bike store for one used dirt bike and one used racing bike?

OpenStudy (anonymous):

there should be options to choose from

OpenStudy (anonymous):

but the way for a solution in such problem is to find the minimum price and maximum price can you do it ?

OpenStudy (anonymous):



OpenStudy (anonymous):

you have to find minimum and maximum prices .. if you can do it then you will know the answer since it will be one of them (minimum or maximum) or in between

OpenStudy (anonymous):

reiu trkje to lazy !

OpenStudy (anonymous):

what would be the minimum price ? 25 + 100 = 125 what would be the maximum ?

OpenStudy (anonymous):

(we already know the answer to the question because there is only one that it is above the minimum

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

maximum is 150

OpenStudy (anonymous):

$91 ?

OpenStudy (anonymous):

no it isnt .. 150 + 75

OpenStudy (anonymous):

look at the question .. we bought one dirt and one race lowest price possible would be 25 + 100 = 125 highest would be 150 + 75 = 225

OpenStudy (anonymous):

so b

OpenStudy (anonymous):

the only possible price is between 125 and 225 ..

OpenStudy (anonymous):

yes .. i hope you understand the way

OpenStudy (anonymous):

for the solution

OpenStudy (anonymous):

no i dont i just guessed but you did help !

OpenStudy (anonymous):

what is the lowest price to pay for a dirt bike ?

OpenStudy (anonymous):

25$

OpenStudy (anonymous):

good and for race is 100 right ?

OpenStudy (anonymous):

so if im buying both race and dirt bikes the lowest price possible is 125

OpenStudy (anonymous):

agree?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so in the same manner .. the highest price possible is 75 + 150 = 225

OpenStudy (anonymous):

so the price that i would pay for a dirt and race bike would range from 125 to 225

OpenStudy (anonymous):

since 125 is minimum and 225 is maximum

OpenStudy (anonymous):

got it now ?